OpenStudy (anonymous):

4 SR 3 + 2 SR 27 -4 SR 18. Solve the radical!

OpenStudy (anonymous):

\[4\sqrt{3} + 2\sqrt{27} - 4\sqrt{18}=4\sqrt{3} + 2\sqrt{3^2\times3} - 4\sqrt{2\times3^2}\] \[=4\sqrt{3} + 6\sqrt{3} - 12\sqrt{2} =10\sqrt{3} - 12\sqrt{2}\]

OpenStudy (anonymous):

yup^